### v2

@@ Appendix C Proof of Section 4
By the law-invariance and monotonicity of $\rho$, we have for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$ and $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, $\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}),$ which implies
@@
Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. Suppose $y<\inf^{\mathbb{Q}_{2}}X-x$. For $z\geqslant x+y$, we have
$\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)$ $=\frac{\mathbb{P}(A_{x}^{c}\cap B_{2})+\mathbb{P}(X\leqslant z,A_{x}\cap B_{2})}{\mathbb{P}(B_{2})}$ (13)
$=1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})},$
where $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$. Moreover, $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)=0$ for $z<x+y$. If $g(\beta)<1$, then there exists $x\in\mathbb{R}$ such that $g(\beta_{x})<1$. By (13), we have $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)\geqslant 1-\beta_{x}$ for all $z\geqslant x+y$, which implies $\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\leqslant x+y\to-\infty$ as $y\to-\infty$ for all $t\in[\beta_{x},1)$. Hence, for $x\in\mathbb{R}$ with $\beta_{x}=\beta$, we have
$\inf_{y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})=\lim_{y\to-\infty}\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)=-\infty.$
Now, we assume $g(\beta)=1$. Direct computation shows for $0<t<\beta_{x}$,
$\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})$ $=\inf\{z\geqslant x+y:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}\geqslant 1-t\}$
@@
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\int_{[0,\beta)}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)$
$=\inf_{x\in\mathbb{R}}\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t)=\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
Let us now consider (ii). If $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then there exists $x\in\mathbb{R}$ such that $\mathbb{P}(B_{1}\cap B_{2}^{c})\geqslant(1-\Lambda(x))\mathbb{P}(B_{1})$. Hence, $A_{x}=\emptyset$. By Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+u(y))=-\infty$.
Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. It follows from Proposition 5 and (C) that
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathbb{E}_{u}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{2}}u((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\left(x+(1-\beta_{x})u(-\infty)+\int_{0}^{\beta_{x}}u\left(\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)-x\right)\mathrm{d}t\right),$
where $u(-\infty)=\lim_{y\to-\infty}u(y)$.
Finally, we consider (iii). Similarly as above, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then by Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+y)=-\infty$.
@@
$=\inf_{x\in\mathbb{R}}\lim_{y\to-\infty}\left\{x+\sup\{z\geqslant y:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant x+z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z)\}\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\to-\infty}\sup\left\{z\geqslant y+x:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:1-\beta_{x}+\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<\beta_{x}-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(X\leqslant z,B_{1}\cap B_{2})}{\mathbb{P}(B_{1}\cap B_{2})}<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}$
@@
$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{B})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B^{c}})\right\}$
$=\inf_{x\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})+u(-\infty)\int_{0}^{\Lambda(x)}\mathrm{VaR}_{t}^{\mathbb{Q}_{1}}(\eta)\mathrm{d}t\right\}.$
Appendix D Proof of Section 5
In this section, we offer all the proofs of the results in Section 5.
@@ Appendix D Proof of Section 5
Next, we show the inverse inequality. For $X_{1}\in\mathcal{X}$, we let $x_{1}=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$. We choose $U_{X_{1}}^{\mathbb{Q}_{1}}$ such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$. This implies $\{F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant x\}=\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant F_{X_{1}}^{\mathbb{Q}_{1}}(x)\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(F_{X_{1}}^{\mathbb{Q}_{1}}(x))\}\subset\{X_{1}\leqslant x\}$ for all $x\in\mathbb{R}$. Hence, $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ for all $w\in\Omega$. By definition, $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant 1-\Lambda(x)$ for all $x\geqslant x_{1}$. This implies $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for $x_{1}\leqslant x\leqslant y$. If $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, then $F_{X_{1}}^{\mathbb{Q}_{1}}(x)=1\geqslant\Lambda_{x_{1},y}(x)$ for $x\geqslant y$. Hence, if $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for all $x\in\mathbb{R}$. Combing all the above results, we have $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant\Lambda_{x_{1},y}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ if $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$. It follows that for $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$,

### v3

@@ Appendix A Proof of all results in Sections 2-5
By the law-invariance and monotonicity of $\rho$, for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, and $y<\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{2}}(X-x)$, we have
$\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}).$
For the case that $\rho^{\mathbb{Q}_{2}}$ is an $\varepsilon$-tail risk measure, let $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-x)$. Then for $z<y$,
$\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)<1-\varepsilon,$
which implies $\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}})\geqslant y$. Combing the above conclusion, we have
$\frac{(\mathbb{Q}_{2}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}\leqslant z)-(1-\varepsilon))_{+}}{\varepsilon}\leqslant\frac{(\mathbb{Q}_{2}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}}\leqslant z)-(1-\varepsilon))_{+}}{\varepsilon},~z\in\mathbb{R}.$
Using the monotonicity of $\rho$ and the fact that $\rho$ is an $\varepsilon$-tail risk measure, we have for $B\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(B)=1-\Lambda(x)$, and $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X-x)$,
$\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{A_{x}}+y\mathds{1}_{A_{x}^{c}})\leqslant\rho^{\mathbb{Q}_{2}}((X-x)\mathds{1}_{B}+y\mathds{1}_{B^{c}}).$
Consequently,
@@
Next, we consider the case $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$. Suppose $y<\mathrm{VaR}_{\varepsilon}^{\mathbb{Q}_{2}}(X)-x$. For $z\geqslant x+y$, we have
$\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)$ $=\frac{\mathbb{P}(A_{x}^{c}\cap B_{2})+\mathbb{P}(\{X\leqslant z\}\cap A_{x}\cap B_{2})}{\mathbb{P}(B_{2})}$ (14)
$=1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})},$
where $\beta_{x}=\frac{(1-\Lambda(x))\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$. Moreover, $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)<\varepsilon$ for $z<x+y$. If $g(\beta)<1$, then there exists $x\in\mathbb{R}$ such that $g(\beta_{x})<1$. By (14), we have $\mathbb{Q}_{2}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}}\leqslant z)\geqslant 1-\beta_{x}$ for all $z\geqslant x+y$, which implies $\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\leqslant x+y\to$-$\infty$ as $y\to$-$\infty$ for all $t\in[\beta_{x},1)$. Hence, for $x\in\mathbb{R}$ with $\beta_{x}=\beta$, we have
$\inf_{y\in\mathbb{R}}\rho_{g}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\leqslant\lim_{y\to-\infty}\int_{0}^{1}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)=-\infty.$
Now, we assume $g(\beta)=1$. Direct computation shows for $0<t<\beta_{x}\wedge\varepsilon$,
$\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})$
$=\inf\left\{z\geqslant x+y:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}\geqslant 1-t\right\}$
@@
$\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\in\mathbb{R}}\int_{[0,\beta_{x}\wedge\varepsilon)}\mathrm{VaR}_{t}^{\mathbb{Q}_{2}}(X\mathds{1}_{A_{x}}+(x+y)\mathds{1}_{A_{x}^{c}})\mathrm{d}g(t)$
$=\inf_{x\in\mathbb{R}}\int_{[0,\beta_{x}\wedge\varepsilon)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta_{x}-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t)=\int_{[0,\beta)}\mathrm{VaR}_{1-\frac{\mathbb{P}(B_{2})(\beta-t)}{\mathbb{P}(B_{1}\cap B_{2})}}^{\mathbb{Q}}(X)\mathrm{d}g(t).$
Let us now consider case (ii). Similarly as above, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then by Proposition 5, we have $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{1}\mathrm{VaR}^{+,\mathbb{Q}_{2}}(X)=\inf_{x,y\in\mathbb{R}}(x+y)=$-$\infty$.
@@
$=\inf_{x\in\mathbb{R}}\lim_{y\to-\infty}\left\{x+\sup\{z\geqslant y:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant x+z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}\leqslant\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z)\}\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\to-\infty}\sup\left\{z\geqslant y+x:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2}\cap\{U_{X}^{\mathbb{P}}<\alpha_{x}\})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:1-\beta_{x}+\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<1-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2})}{\mathbb{P}(B_{2})}<\beta_{x}-\Lambda_{1}(z-x)\right\}$
$=\inf_{x\in\mathbb{R}}\sup\left\{z\in\mathbb{R}:\frac{\mathbb{P}(\{X\leqslant z\}\cap B_{1}\cap B_{2})}{\mathbb{P}(B_{1}\cap B_{2})}<1-\frac{\mathbb{P}(B_{1})\Lambda(x)+\mathbb{P}(B_{2})\Lambda_{1}(z-x)}{\mathbb{P}(B_{1}\cap B_{2})}\right\}$
@@
$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\inf_{t\in\mathbb{R}}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}((X-t)_{+}\mathds{1}_{B})\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{t\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{t+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{1}}(\eta(X-t)_{+}\mathds{1}_{B})+\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}((X-t)_{+}\mathds{1}_{B\cap N})\right\}$
@@
$=\inf_{x\in\mathbb{R}}\inf_{\mathbb{Q}_{1}(B)=1-\Lambda(x)}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{B})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B^{c}})\right$.
$~~~~\left.+\mathbb{E}^{\mathbb{Q}_{2}}(u(X-x)\mathds{1}_{B\cap N})+u(-\infty)\mathbb{E}^{\mathbb{Q}_{2}}(\mathds{1}_{B^{c}\cap N})\right\}$
$=\inf_{x\in\mathbb{R}}\left\{x+\mathbb{E}^{\mathbb{Q}_{1}}(\eta u(X-x)\mathds{1}_{\{U_{\eta}^{\mathbb{Q}_{1}}<1-\Lambda(x)\}})+u(-\infty)\left(\mathbb{Q}_{2}(N)+\int_{0}^{\Lambda(x)}\mathrm{VaR}_{t}^{\mathbb{Q}_{1}}(\eta)\mathrm{d}t\right)\right\}.$
@@
Next, we show the inverse inequality. For $X_{1}\in\mathcal{X}$, we let $x_{1}=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$. We choose $U_{X_{1}}^{\mathbb{Q}_{1}}$ such that $\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant t\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(t)\}$ for all $t\in(0,1)$. This implies $\{F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant x\}=\{U_{X_{1}}^{\mathbb{Q}_{1}}\leqslant F_{X_{1}}^{\mathbb{Q}_{1}}(x)\}\subset\{X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(F_{X_{1}}^{\mathbb{Q}_{1}}(x))\}\subset\{X_{1}\leqslant x\}$ for all $x\in\mathbb{R}$. Hence, $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ for all $w\in\Omega$. By definition, $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant 1-\Lambda(x)$ for all $x\geqslant x_{1}$. This implies $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for $x_{1}\leqslant x\leqslant y$. Suppose $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}<\infty$. If $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, then $F_{X_{1}}^{\mathbb{Q}_{1}}(x)=1\geqslant\Lambda_{x_{1},y}(x)$ for $x\geqslant y$. Hence, if $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$, we have $F_{X_{1}}^{\mathbb{Q}_{1}}(x)\geqslant\Lambda_{x_{1},y}(x)$ for all $x\in\mathbb{R}$. Combing all the above results, we have $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant\Lambda_{x_{1},y}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ if $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$. It follows that for $y\geqslant\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$,
@@
If $\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}=\infty$, then let $X_{m}=X_{1}\wedge m$ for $m\geqslant 2$. Using the above conclusion, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{m})+\rho^{\mathbb{Q}_{2}}(X-X_{m})\geqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
Direct computation gives $\lim_{m\to\infty}\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{m})=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$. Note that $X-X_{m}\downarrow X-X_{1}$ as $m\to\infty$. Using the fact that $\rho$ is continuous from above, we have $\lim_{m\to\infty}\rho^{\mathbb{Q}_{2}}(X-X_{m})=\rho^{\mathbb{Q}_{2}}(X-X_{1})$. Hence, letting $m\to\infty$, we obtain
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})\geqslant\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$